# Optimal Pure Differentially Private Sparse Histograms in Near-Linear Deterministic Time

### v1

@@ Appendix A Probabilities Inequalities
$\Pr\left[\lvert{X}\rvert\geq t\right]=\frac{2\cdot e^{-\varepsilon\cdot t}}{1+e^{-\varepsilon}}.$ (14)
Proof of A.2.
@@
$\Pr\left[\lvert{X}\rvert\geq t\right]$ $=2\cdot\frac{1-e^{-\varepsilon}}{1+e^{-\varepsilon}}\cdot\sum_{i=t}^{\infty}e^{-\varepsilon\cdot i}=2\cdot\frac{1-e^{-\varepsilon}}{1+e^{-\varepsilon}}\cdot\frac{e^{-\varepsilon\cdot t}}{1-e^{-\varepsilon}}=2\cdot\frac{e^{-\varepsilon\cdot t}}{1+e^{-\varepsilon}}.$ (15)
@@
[…]
@@
$f(0)\doteq\lim_{t\downarrow 0}f(t)\in(0,\infty],\quad f^{*}(0)\doteq\lim_{t\rightarrow\infty}\frac{f(t)}{t},\quad f(0)\doteq 0\doteq 0,\quad f^{*}(0)\doteq 0\doteq 0.$ (17)
@@
$(K\mu)(B)\doteq\int_{\mathcal{X}}K(B\mid x)\,\mu(dx),\quad\forall B\in\mathcal{B}.$ (18)
Proposition A.1 (Data Processing Inequality (see, e.g., Theorem 14 of Liese and Vajda (2006))).
@@
$\mathds{D}_{f}\left({K\mu},\,{K\nu}\right)\leq\mathds{D}_{f}\left({\mu},\,{\nu}\right).$ (19)
Proposition A.2 (Sendler (1975), Lemma 2.1; Steerneman (1983), Theorem 3.1).
@@
$\mathds{TV}({\times_{k\in[n]}\mu_{k}},\,{\times_{k\in[n]}\nu_{k}})\leq\sum_{k\in[n]}\mathds{TV}({\mu_{k}},\,{\nu_{k}}),$ (20)
where ×$\times_{k\in[n]}\mu_{k}$ and ×$\times_{k\in[n]}\nu_{k}$ are the product measures on ${\left({\Omega^{n},\mathcal{F}^{n}}\right)}$ constructed from $\mu_{k}$ and $\nu_{k}$, respectively.
@@
$\mathds{TV}({X},\,{Y})\leq\sum_{k\in[n]}\mathds{TV}({X_{k}},\,{Y_{k}}).$ (21)
Appendix B Alias Method Initialization
In this section, we describe how to initialize the two arrays $\vec{a}$ and $\vec{b}$ used in the Alias method.
@@ Appendix B Alias Method Initialization
3:$\bar{m}\leftarrow 2^{{\lceil{\log_{2}m}\rceil}}$triangleright rounding $m$ up to the nearest power of $2$
@@
    $\vec{c}[k]+\sum_{i\in\bar{m}}{\left({\vec{b}[i]\cdot\mathds{1}_{\left[\vec{a}[i].\textsc{first}=k\right]}+(1-\vec{b}[i])\cdot\mathds{1}_{\left[\vec{a}[i].\textsc{second}=k\right]}}\right)}=p_{k}\cdot\bar{m}.$ (22)

### v2

@@ A.1. Probabilities Inequalities
(5) $\Pr\left[\lvert{X}\rvert\geq t\right]=\frac{2\cdot e^{-\varepsilon\cdot t}}{1+e^{-\varepsilon}}.$
Proof of ˜A.2.
@@
(6) $\Pr\left[\lvert{X}\rvert\geq t\right]$ $=2\cdot\frac{1-e^{-\varepsilon}}{1+e^{-\varepsilon}}\cdot\sum_{i=t}^{\infty}e^{-\varepsilon\cdot i}=2\cdot\frac{1-e^{-\varepsilon}}{1+e^{-\varepsilon}}\cdot\frac{e^{-\varepsilon\cdot t}}{1-e^{-\varepsilon}}=2\cdot\frac{e^{-\varepsilon\cdot t}}{1+e^{-\varepsilon}}.$
@@
(7) $\mathds{D}_{f}\left({\mu},\,{\nu}\right)=\int_{pq>0}f{\left({\frac{p}{q}}\right)}\,d\nu+f(0)\cdot\nu(p=0)+f^{*}(0)\cdot\mu(q=0),$
@@
(8) $f(0)\doteq\lim_{t\downarrow 0}f(t)\in(0,\infty],\quad f^{*}(0)\doteq\lim_{t\rightarrow\infty}\frac{f(t)}{t},\quad f(0)\doteq 0\doteq 0,\quad f^{*}(0)\doteq 0\doteq 0.$
@@
(9) $(K\mu)(B)\doteq\int_{\mathcal{X}}K(B\mid x)\,\mu(dx),\quad\forall B\in\mathcal{B}.$
Proposition A.0 (Data Processing Inequality (see, e.g., Theorem 14 of (Liese and Vajda, 2006))).
@@
(10) $\mathds{D}_{f}\left({K\mu},\,{K\nu}\right)\leq\mathds{D}_{f}\left({\mu},\,{\nu}\right).$
Proposition A.0 ((Sendler, 1975), Lemma 2.1; (Steerneman, 1983), Theorem 3.1).
@@
(11) $\mathds{TV}({\times_{k\in[n]}\mu_{k}},\,{\times_{k\in[n]}\nu_{k}})\leq\sum_{k\in[n]}\mathds{TV}({\mu_{k}},\,{\nu_{k}}),$
where $\times_{k\in[n]}\mu_{k}$ and $\times_{k\in[n]}\nu_{k}$ are the product measures on $\left({\Omega^{n},\mathcal{F}^{n}}\right)$} constructed from $\mu_{k}$ and $\nu_{k}$, respectively.
@@
(12) $\mathds{TV}({X},\,{Y})\leq\sum_{k\in[n]}\mathds{TV}({X_{k}},\,{Y_{k}}).$
A.2. The Poisson Approximation
When throwing $m$ balls into $d$ bins independently and uniformly at random, the load of any fixed bin follows the binomial distribution $\operatorname{\mathbb{B}in}({m},{1/d})$. In the regime where $m/d$ is fixed and $m\to\infty$, this binomial distribution converges to a Poisson distribution.
Definition A.8 (Poisson Distribution).
A discrete Poisson random variable $X$ with parameter $\lambda>0$ has probability mass function
$\Pr\left[X=j\right]=e^{-\lambda}\frac{\lambda^{j}}{j!},\qquad j\in\mathbb{N}.$
More generally, in many settings it is fruitful to approximate the joint distribution of the loads of all bins by assuming that each bin receives an independent Poisson number of balls with mean $m/d$.
Theorem A.9 (Poisson Approximation Theorem (Mitzenmacher and Upfal, 2005)).
Suppose $m$ balls are thrown into $d$ bins independently and uniformly at random, and let $X_{i}$ denote the number of balls in bin $i$, for $1\leq i\leq d$. Let $Y_{1},\dots,Y_{d}$ be independent Poisson random variables, each with mean $m/d$. For any nonnegative function $f(x_{1},\ldots,x_{d})$,
$\operatorname*{\mathbb{E}}\left[f(X_{1},\ldots,X_{d})\right]\;\leq\;e\sqrt{m}\,\operatorname*{\mathbb{E}}\left[f(Y_{1},\ldots,Y_{d})\right].$
In particular, for any event $\mathcal{E}$, taking $f$ to be its indicator function yields
[…]
We also use the following standard tail bounds for Poisson random variables.
Theorem A.10 (Poisson Tail Bounds (Mitzenmacher and Upfal, 2005)).
Let $Y\sim\operatorname{\mathbb{P}oisson}\left({\mu}\right)$. For any $0\leq\alpha\leq 1$,
$\Pr\left[Y\leq\mu(1-\alpha)\right]\;\leq\;\exp\!\left(-\frac{\alpha^{2}\mu}{2}\right),\qquad\Pr\left[Y\geq\mu(1+\alpha)\right]\;\leq\;\exp\!\left(-\frac{\alpha^{2}\mu}{3}\right).$
A.3. Alias Method
The Alias method (Walker, 1977), denoted by $\mathcal{M}_{\mathbb{A}lias}$}, is an efficient array-based sampling algorithm that, given a distribution $\mu=\{{p_{0},\ldots,p_{m-1}}\}$ over $\mathbb{N}_{<{m}}$, has $O(m)$ initialization time and memory usage, and then draw samples from $\mu$ in $O(1)$ time. The pseudocode is given in Algorithm˜3. During the initialization phase, the algorithm constructs two arrays $\vec{a}$ and $\vec{b}$ of length $\bar{m}\geq m$. The array $\vec{a}\in\mathbb{N}_{<{m}}^{\bar{m}}$ is called the alias array, and $\vec{b}\in[0,1]^{\bar{m}}$ stores Bernoulli probabilities.
Algorithm 3 Alias Method $\mathcal{M}_{\mathbb{A}lias}$}
1:Procedure: Initialization
2: Input: $m\in\mathbb{N}^{+}$, $p_{0},\ldots,p_{m-1}\in{({0,1})}$
3:Construct arrays $\vec{a}$ and $\vec{b}$ as in Proposition˜A.11 (Walker, 1977);
4:Procedure: Sample
5: Output: Random variable $Z\in\mathbb{N}_{<{m}}$
[…]
7:if $B=1$ then return $Z\leftarrow I$
8:if $B=0$ then return $Z\leftarrow\vec{a}[I]$
Sampling Procedure ($\mathcal{M}_{\mathbb{A}lias}.\textsc{\small Sample}$}):
To sample from $\mu$, the algorithm proceeds as follows: (1) sample a uniform random index $I\in\mathbb{N}_{<{\bar{m}}}$, (2) sample a Bernoulli random variable $B$ that equals $1$ probability $\vec{b}[I]$, (3) if $B=1$, return $I$; otherwise, return $\vec{a}[I]$. That is why $\vec{a}$ is called the alias array: it specifies the alternative index to return when the Bernoulli test fails.
Proposition A.0 ((Walker, 1977)).
For completeness, we also include the pseudocode as well as the proof of the construction of $\vec{a}$ and $\vec{b}$. It is worth noting that Walker (1977) assumes real arithmetic, which is infeasible in the Word-RAM model used in this paper. In Section˜B.3, we relax this assumption and apply the Alias method to approximately sample from a discrete distribution.
Alias Method Initialization
We now describe how to initialize the two arrays $\vec{a}$ and $\vec{b}$ used in the Alias method.
@@ A.3. Alias Method
[…]
@@
    (13) $\vec{c}[k]+\sum_{i\in\bar{m}}{\left({\vec{b}[i]\cdot\mathds{1}_{\left[\vec{a}[i].\textsc{first}=k\right]}+(1-\vec{b}[i])\cdot\mathds{1}_{\left[\vec{a}[i].\textsc{second}=k\right]}}\right)}=p_{k}\cdot\bar{m}.$